Assignment: Capital Budgeting Decisions

Your company is considering undertaking a project to expand an existing product line. The required rate of return on the project is 8% and the maximum allowable payback period is 3 years.

time

0

1

2

3

4

5

6

Cash flow

$ 10,000

2,400

4,800

3,200

3,200

2,800

2,400

Questions

Evaluate the project using each of the following methods. For each method, should the project be accepted or rejected? Justify your answer based on the method used to evaluate the project’s cash flows.

  1. Payback period
  2. Internal Rate of Return (IRR)
  3. Simple Rate of Return
  4. Net Present Value
Answer #1

Total investment in Period 0 = $10000

Total returns in 3 years = 2400 + 4800 + 3200 = $10400

Considering a simple return method, the return in 3 years is greater than investment hence project is accepted.

The present value of 3 years return (See table below) = 2222 + 4115 + 2540 = $8877

Considering present value of cash-flows, the project is rejected as per the payback method as Present value of 3 years inflow is lower than initial outflow.

IRR = 22.69%.  Project is accepted as IRR is greater than the discount rate. IRR is the rate at which NPV becomes zero. (See image for calculation, cell B8)

Formulas:

Total return in Year (1-6) = 2400 + 4800+3200+3200+2800+2400 = $18800

Rate of return = Total returns / Capital outflow = 18800 / 10000 = 188% , The project should be accepted.

Discount rate (R%) 8%
Time = n 0 1 2 3 4 5 6 Total
Cash flow ($10,000) 2,400 4,800 3,200 3,200 2,800 2,400
PV Factor = 1/ (1+r)^n 1 0.9259 0.8573 0.7938 0.7350 0.6806 0.6302
Present value of Cash flow = Cash Flow * PV factor ($10,000) $2,222 $4,115 $2,540 $2,352 $1,906 $1,512 $4,648
NPV $4,648
IRR 22.69%

With the NPV method the project is accepted as NPV is positive.